# machine learning, optical communications, passive optical network, equalization, optical performance monitoring, modulation format identification, bit error rate, optical signal-to-noise ratio, nonlinearities

### v1

@@ 7 Conclusion
In this survey, we have undertaken a comprehensive examination of powerful machine learning models that exhibit the potential to achieve robust equalization in cost-sensitive short-reach systems, with a particular focus on PONs. Our objective has been to explore these models’ capacity to operate efficiently and deliver effective computational performance. For the first time, we have classified the current models into three distinct types and conducted an extensive analysis of their core concepts, highlighting their differences, similarities, and the underlying insights they provide. Additionally, we have presented a simplified complexity analysis considering various input sizes. In the final stages of our survey, we have also investigated the potential of machine learning solutions in addressing the challenges associated with hardware implementation and complexity. We firmly believe that this survey will serve as a valuable resource, inspiring future research endeavors to develop efficient models explicitly tailored for short-reach and PON systems.

### v2

@@ 7 Conclusion
In this survey, we have undertaken a comprehensive examination of powerful machine learning models that exhibit the potential to achieve robust equalization in cost-sensitive short-reach optical systems, with a particular focus on PONs. Our objective has been to explore these models’ capacity to operate efficiently and deliver effective computational performance. For the first time, we have classified the current models into three distinct types and conducted an extensive analysis of their core concepts, highlighting their differences, similarities, and the underlying insights they provide. Additionally, we have presented a simplified complexity analysis considering various input sizes. In the final stages of our survey, we have also investigated the potential of machine learning solutions in addressing the challenges associated with hardware implementation and complexity. We firmly believe that this survey will serve as a valuable resource, inspiring future research endeavors to develop efficient models explicitly tailored for short-reach and PON systems.